# Revision of two-temperature magnetically arrested flows onto a black hole

### v1

@@ Abstract
We revisit radiative properties of 3-D GRMHD two-temperature magnetically arrested disk (MAD) models in which electrons are heated up by magnetic turbulent cascade. We focus on studying the models emission which characteristics include variability in both total intensity and linear/circular polarizations as well as rotation measures at energies around synchrotron emission peak in millimeter waves. We find that radiative properties of MAD models with turbulent electron heating are well converged with respect to the numerical grid resolution, which has not been demonstrated before. We compare radiation from the two-temperature simulations with turbulent heating to single-temperature models with electron temperatures calculated based on commonly used $R(\beta)$ prescription. We find that the two-temperature models do not significantly overperform the $R(\beta)$ models and in practice may be indistinguishable from the $R(\beta)$ models. Accounting for physical effects such as radiative cooling and non-thermal electron distribution function makes a weak impact on properties of millimeter emission. Models are scaled to Sgr A* - an accreting black hole in the center of our galaxy, and compared to the most complete observational datasets. We point out consistencies and inconsistencies between the MAD models and observations of this source and discuss future prospects for GRMHD simulations.
@@ 1 Introduction
Magnetically arrested disks (hereafter MADs) is a class of solutions of black hole accretion initially developed for radiatively inefficient accretion flows (RIAF) (Igumenshchev et al., 2003; Narayan et al., 2003; Proga & Begelman, 2003; Tchekhovskoy et al., 2011; McKinney et al., 2012; Narayan et al., 2022). In MADs realized in general relativistic magnetohydrodynamics (GRMHD) simulations, the magnetic field flux accumulated near the black hole event horizon is extreme, which has two important consequences. First, the power of jets produced in MAD simulations is significant, making the jet opening angle rather large ($\sim 55$ deg, Chael et al. 2019). This is consistent with what is observed in sub-Eddington accreting black hole systems such as M87 (Walker et al., 2018; Kim et al., 2018). MADs near black hole horizon are further supported in M87 by the first polarimetric images from the Event Horizon Telescope (EHT; Event Horizon Telescope Collaboration et al. 2021). Second, the jet magnetosphere in MADs is often reconnecting at the equatorial plane of the simulation causing magnetic field flux eruptions into the disk body. Such flux eruptions has been proposed to explain flaring behaviour of another sub-Eddington accreting supermassive black hole, Sgr A* (Dexter et al., 2020b; Porth et al., 2021; Wielgus et al., 2022b). It has been also recently found that horizon scale polarimetric images of Sgr A* are more consistent with MADs (Event Horizon Telescope Collaboration et al., 2024). While less magnetized standard and normal evolution accretion disks (SANEs) are not completely ruled out by observations of any of the sources, MAD RIAF solution become a leading candidate to explain observations of two best-resolved aforementioned accreting supermassive black holes.
@@
One of the most uncertain parts of MADs (but also SANEs) is how they generate emission. In the aforementioned class of systems, where the accreting plasma is collisionless, electron and ion distribution functions remain unknown. Typically, to mimic the collisionless effects in models we assume that the plasma has two-temperature structure and we “paint” electron temperatures on the top of the GRMHD models according to some parameterized law. One example of such parameterized law is the $R(\beta$) prescription of Mościbrodzka et al. 2016 used in Event Horizon Telescope Collaboration et al. 2022, where model parameter $R(\beta)\equiv T_{\rm i}/T_{\rm e}(\beta)$ is a ratio of non-emitting ion to emitting electron temperatures that depends on the local plasma $\beta(\equiv P_{\rm mag}/P_{\rm gas}$) parameter. Other than the $R(\beta)$ models has been proposed (Mościbrodzka & Falcke, 2013; Mościbrodzka et al., 2014; Chan et al., 2015; Gold et al., 2017; Anantua et al., 2020). Alternatively, one can assume a non-thermal electron distribution function, but those are typically also parametric (see e.g., Davelaar et al. 2018; Scepi et al. 2022; Cruz-Osorio et al. 2022; Fromm et al. 2022; Zhao et al. 2023 for a recent work and Chael et al. 2017; Petersen & Gammie 2020 for examples of models with evolution of non-thermal electrons). Finally one can model two-temperature GRMHD flows where electron temperatures are followed by a separate equation which includes non-adiabatic (viscous), sub-grid heating terms based on plasma or particle-in-cell considerations (Howes 2010; Rowan et al. 2017; Kawazura et al. 2019) and radiative cooling terms (e.g., Ryan et al. 2018).
@@
The method to track electron temperatures in GRMHD simulations assuming a sub-grid model for ion and electron heating has been introduced by Ressler et al. (2015). Sadowski et al. (2017) further developed the idea to add the ion/electron pressure as well as self-consistent variable adiabatic index to the GRMHD evolution. Two-temperature 2-D and 3-D GRMHD models have been carried out in the past, specifically for Sgr A* (Ressler et al. 2017; Sadowski et al. 2017; Chael et al. 2018; Dexter et al. 2020a; Jiang et al. 2023) and M87 (Ryan et al. 2018; Mizuno et al. 2021; Dihingia et al. 2023).
@@
In this paper we revisit the two-temperature GRMHD MAD simulations. The mentioned previous studies use several numerical codes assuming various: torus sizes, magnetic field topologies, adiabatic indices and grid resolutions to investigate the problem. It was not clear to us that the radiative output of these models is converged for MAD models with setups comparable to those in the EHT simulation library (Event Horizon Telescope Collaboration et al., 2022). The second motivation for the revision of two-temperature simulations is the fact that the MAD models with $R(\beta)$ prescription have typically too variable total intensity when compared to the observations of Sgr A* (Wielgus et al., 2022a; Event Horizon Telescope Collaboration et al., 2022). The origin of the discrepancy between the models and observations is unknown. One possibility is that the $R(\beta)$ model, which was initially developed for the SANE simulations, is not appropriate for MADs and more sophisticated electron model should be adopted. One could also ask, how physically motivated is the $R(\beta)$ model for electron temperatures when considering MADs. The latter is relevant for all sub-Eddington accreting black hole systems. Finally, the measurement of black hole spin in EHT and other sources is typically accretion model dependent. A robust understanding of the dissipation processes in accreting plasma near the event horizon is critical for reliable spin estimates.
@@
In this work, we investigate how numerical model parameters impact radiative (total intensity and polarimetric) characteristics of the two-temperature MAD simulations. We first demonstrate that they are independent of grid resolution used in GRMHD simulations. Next, we evolve models for longer times and compare the two-temperature models to $R(\beta)$ models. Finally, having checked that the results are weakly dependent on the exact shape of the electron distribution function and that they remain unchanged even when radiative cooling is introduced (scaling models to Sgr A* system), we compare the prograde and retrograde two-temperature MAD models to the newest multi-frequency polarimetric millimeter observational data of Sgr A*. In contrast to most of the previous two-temperature simulations, in our comparisons we focus on variability and polarization, which together are significantly more informative than the time-averaged or total intensity comparisons. We discuss the future prospects for investigating thermodynamics of the GRMHD simulations.
@@
The paper is structured as follows. In Section 2 we describe our 3-D two-temperature GRMHD simulations setups and outline the details of radiative transfer modeling. In Section 3, we report the results and compare the models to selected observations of Sgr A*. We discuss the results and conclude in Section 4.
@@ 2.1 GRMHD simulations
MAD simulations are carried by means of ebhlight, general relativistic radiation magnetohydrodynamics code developed and made public by Ryan et al. (2015). ebhlight is a relativistic, second-order, conservative, constrained-transport code for stationary spacetimes. While the fluid part of the code is based on harm routines (Gammie et al., 2003), the radiation part is based on grmonty scheme (Dolence et al., 2009).
All simulations are set in geometrized units where the lengthscale and timescale are set by the black hole mass: ${\mathcal{L}}\equiv GM/c^{2}$ and ${\mathcal{T}}\equiv GM/c^{3}$ (assuming $G=c=1$ both time and distances are measured in units of mass, M). The simulations start from the Fishbone-Moncrief torus (Fishbone & Moncrief, 1976) at the equatorial plane of a Schwarzschild or Kerr black hole. The torus is described by two parameters: inner radius $r_{\rm in}=20M$ and radius of the pressure maximum, $r_{\rm max}=41M$. The torus is seeded with weak poloidal magnetic fields described by vector potential:
@@
where $r,\theta$ are the radius and polar angle in Kerr-Schild coordinates, parameter $r_{0}=400$, $\rho$ is the plasma density. The initial magnetic field is renormalized so that initially plasma $\beta_{\rm max}=100$.
The simulations use ideal equation of state with constant adiabatic index of $\gamma_{ad}=13/9$, i.e., we assume that plasma is a pure hydrogen with non-relativistic protons and relativistic electrons, for which adiabatic indices are $\gamma_{\rm p}=5/3$ and $\gamma_{\rm e}=4/3$, respectively.
All simulations are carried out in mixed modified Kerr-Schield logarithmic coordinates where resolution is focused on equatorial plane and towards the central region close to the black hole horizon. The grid stretches from within the black hole event horizon until $r_{\rm out}=1,000$M.
The grid refinement convergence test runs are integrated for $10,000$M assuming four grid resolutions listed in Table 1. Convergence runs are done for black hole spin $a_{*}=0$. Later in the paper we consider both prograde and retrograde fiducial models (also listed in Table 1) with the following spin values $a_{*}$=-0.9375,-0.5,0,0.5,0.9375. Prograde and zero spin models are evolved until $30,000$M, and retrograde models are evolved for a shorter times, only until $14,000$M. Fiducial models have grid resolutions ($N_{r},N_{\theta},N_{\phi}$)=(240,120,128) (for $a_{*}$=-0.5,0,0.5) and ($N_{r},N_{\theta},N_{\phi}$)=(266,120,128) (for $a_{*}$=-0.9375,0.9375). All models above are run with radiative modules of ebhlight turned off. We carry out one test run with radiative effects turned on. The parameters of the exploratory GRRMHD model are shown in Table 1.
@@ 2.2 Radiative transfer
[…]
@@ 2.3 Electron distribution functions
Following scheme of Ressler et al. (2015), the ebhlight code tracks total and electron entropies that are used to calculate the non-adiabatic (viscous) gas heating rates and evolve electron temperatures. The scheme requires providing a sub-grid model for proton-to-electron heating ratio, $Q^{+}_{\rm p}/Q^{+}_{\rm e}$. In this work we adopt prescription for $Q^{+}_{\rm p}/Q^{+}_{\rm e}$ developed by Kawazura et al. (2019) (henceforth model K) which approximates dissipation in a turbulent cascade with a functional form:
@@
In this model most of the dissipation goes to protons in high $\beta$ plasma regions, while electrons receive most of the heating in low $\beta$ regions. Qualitatively, model K is very similar to turbulent heating model by Howes (2010). Other $Q^{+}_{\rm p}/Q^{+}_{\rm e}$ options are discussed in Section 4.
In ideal GRMHD models studied here using conservative code, the total viscous heating is produced by truncation errors at the numerical grid level. One may conclude that no matter what fraction of this heating goes to the electrons, this heating will be completely artificial. However, for the turbulent torus problem, the grid-scale dissipation is set by the large-scale turbulence in the problem (see section 3.1 in Ressler et al. 2015 for more detailed discussion). The numerical scheme for electron heating implemented in ebhlight has been carefully tested against some analytic problems by Ressler et al. (2015) and Ryan et al. (2017) as well as Sadowski et al. (2017) who, except for the mentioned modifications, follows the same idea when evolving electron temperatures alone. We are not going to repeat these tests here. In ebhlight the electron evolution also does not include Coulomb couplings, unless radiative transfer is activated (see Section 3.2.1). Those are however not important in low density gas considered here. In ebhlight the electron temperature tracking is passive; i.e. the electron pressure is not accounted for in the GRMHD equations. More importantly, the proton (non-adiabatic/viscous) heating is not taken into account in the GRMHD equations as well; instead the gas internal energy is evaluated from the total energy using an inversion scheme implemented in the code assuming $\gamma_{ad}$. Proton temperatures, if needed, are computed using the total internal energy $u$, plasma density $\rho$ and $\gamma_{\rm p}=5/3$ ($\Theta_{\rm p}\equiv(\gamma_{\rm p}-1)u/\rho$).
The simulations are two-temperature but to make a connection with the previous studies we use the same GRMHD runs and compute images/light-curves using parametric electron temperature model $R(\beta)$. In this model proton temperatures are calculated from GRMHD quantities and electron temperature is found using formula below:
@@
where parameter $R_{\rm low}$ and $R_{\rm high}$ are temperature ratios that describe the proton-to-electron temperature ratio in strongly (low plasma $\beta$) and weakly (high plasma $\beta$) magnetized regions, respectively. To be consistent with Event Horizon Telescope Collaboration et al. (2022, 2024) we calculate light-curves for $R_{\rm high}=1,10,40,160$ and $R_{\rm low}=1$. The dimensionless electron temperature that is passed to the synchrotron transfer coefficients in ipole is calculated from:
@@
with internal energy $u$ and rest-mass density $\rho$ provided by GRMHD model. The electron model assumes $\gamma_{\rm e}=4/3$ and $\gamma_{\rm p}=5/3$, self-consistent with GRMHD simulations in which electrons are relativistic and protons are non-relativistic.
Models with Kawazura et al. (2019) prescription for $Q^{+}_{\rm p}/Q^{+}_{\rm e}$ are tagged K and models with parametric electron temperatures are labelled R1-R160 (e.g., model with $R_{\rm low}=1$ and $R_{\rm high}=160$ is denoted R160). The latter models are referred to as $R(\beta)$ models.
It is reasonable to assume that the energy from the turbulent heating is used to form a non-thermal rather than purely thermal electron distribution function. We consider an exploratory non-thermal model using $\kappa(w,\kappa)$ distribution function (Xiao, 2006) where the distribution parameter
$w=\frac{(\kappa-3)}{\kappa}\Theta_{e}.$ (5)
Here $\Theta_{e}$ is calculated form the electron entropy in exactly same manner as in purely thermal K models (see Ressler et al. 2015). Equation 5 states that the entire turbulent energy is distributed into $\kappa$ function. The $\kappa$ parameter is globally constant. Inspired by the Solar wind studies we assume $\kappa=4.25$ (e.g., Livadiotis et al., 2018). In ray-tracing, the radiative transfer coefficients for the $\kappa$ distribution function are adopted from Mościbrodzka & Gammie (2024). No cutoff is applied to the $\kappa$ distribution function at high energies because it makes negligible difference when modeling (sub-)millimeter emission.
All radiative transfer simulations are carried out only in regions where $\sigma<\sigma_{\rm cut}=1$ (where magnetization parameter $\sigma\equiv 2P_{\rm mag}/\rho c^{2}$). This is a standard procedure to avoid modeling emission from near-vacuum jet regions in GRMHD model where proton and electron temperatures are inaccurate (e.g., Event Horizon Telescope Collaboration et al., 2022).

### v3

@@ Abstract
We revisit the radiative properties of 3D general relativistic magnetohydrodynamics (GRMHD) two-temperature magnetically arrested disk (MAD) models in which electrons are heated by a magnetic turbulent cascade. We focus on studying the model emission, whose characteristics include variability in both total intensity and linear/circular polarizations as well as rotation measures at energies around the synchrotron emission peak in millimeter waves. We find that the radiative properties of MAD models with turbulent electron heating are well converged with respect to the numerical grid resolution, which has not been demonstrated before. We compare radiation from two-temperature simulations with turbulent heating to single-temperature models with electron temperatures calculated based on the commonly used $R~{}(\beta)$ prescription. We find that the self-consisitent two-temperature models with turbulent heating do not significantly outperform the $R~{}(\beta)$ models and, in practice, may be indistinguishable from the $R~{}(\beta)$ models. Accounting for physical effects such as radiative cooling and the nonthermal electron distribution function makes a weak impact on properties of millimeter emission. Models are scaled to Sgr A*, an accreting black hole in the center of our galaxy, and compared to the most complete observational datasets. We point out the consistencies and inconsistencies between the MAD models and observations of this source and discuss future prospects for GRMHD simulations.
@@ 1 Introduction
Magnetically arrested disks (hereafter MADs) are a class of black hole accretion solutions initially developed for radiatively inefficient accretion flows (RIAFs; Igumenshchev et al. 2003; Narayan et al. 2003, 2022; Proga & Begelman 2003; Tchekhovskoy et al. 2011; McKinney et al. 2012). In MADs realized in general relativistic magnetohydrodynamics (GRMHD) simulations, the magnetic field flux accumulated near the black hole event horizon is extreme, which has two important consequences. First, the power of jets produced in MAD simulations is significant, which makes the jet opening angle rather large ($\sim 55^{\circ}$; Chael et al. 2019). This is consistent with what is observed in sub-Eddington accreting black hole systems such as M87 (Kim et al., 2018; Walker et al., 2018). MADs near the black hole horizon are further supported in M87 by the first polarimetric images from the Event Horizon Telescope (EHT; Event Horizon Telescope Collaboration et al. 2021). Second, the jet magnetosphere in MADs is often reconnecting at the equatorial plane of the simulation causing magnetic field flux eruptions into the disk body. Such flux eruptions have been proposed to explain the flaring behavior of another sub-Eddington accreting supermassive black hole, Sgr A* (Dexter et al., 2020b; Porth et al., 2021; Wielgus et al., 2022b). It has also been recently found that horizon-scale polarimetric images of Sgr A* are more consistent with MADs (Event Horizon Telescope Collaboration et al., 2024). Although less magnetized standard and normal evolution accretion disks (SANEs), or wind-fed accretion or tilted-disk models (e.g., Ressler et al. 2019; Chatterjee et al. 2020), are not completely ruled out by observations of any of the sources, the MAD RIAF solution became a leading candidate to explain the observations of the two best-resolved aforementioned accreting supermassive black holes.
@@
One of the most uncertain parts of MADs (but also SANEs) is how they generate emission. In the aforementioned class of systems, where the accreting plasma is collisionless, electron and ion distribution functions remain unknown. Typically, to mimic the collisionless effects in models, we assume that the plasma has two-temperature structure, and we “paint” electron temperatures on the top of the GRMHD models according to some parameterized law. An example of such a parameterized law for the thermal distribution of electrons is the $R~{}(\beta$) prescription of Mościbrodzka et al. 2016 used in Event Horizon Telescope Collaboration et al. 2022, where the model parameter $R~{}(\beta)\equiv T_{\rm i}/T_{\rm e}(\beta)$ is a ratio of the nonemitting ion to the emitting electron temperatures that depends on the local plasma $\beta(\equiv P_{\rm mag}/P_{\rm gas}$) parameter. Other than the $R~{}(\beta)$ models has been proposed by Mościbrodzka & Falcke (2013); Mościbrodzka et al. (2014); Chan et al. (2015); Gold et al. (2017); Anantua et al. (2020). Alternatively, one can assume a nonthermal electron distribution function, but those are typically also parametric (see, e.g., Davelaar et al. 2018; Cruz-Osorio et al. 2022; Fromm et al. 2022; Scepi et al. 2022; Zhao et al. 2023 for recent work and Chael et al. 2017; Petersen & Gammie 2020 for examples of models with evolution of nonthermal electrons). Finally, one can model two-temperature GRMHD flows where electron temperatures are followed by a separate equation that includes nonadiabatic (viscous), subgrid heating terms based on plasma or particle-in-cell considerations (Howes 2010; Rowan et al. 2017; Kawazura et al. 2019) and radiative cooling terms (e.g., Ryan et al. 2018).
@@
The method of tracking electron temperatures in GRMHD simulations assuming a subgrid model for ion and electron heating has been introduced by Ressler et al. (2015). Sadowski et al. (2017) further developed the idea to add the ion/electron pressure as well as a self-consistent variable adiabatic index to the GRMHD evolution. Two-temperature 2D and 3D GRMHD models have been carried out in the past, specifically for Sgr A* (Ressler et al. 2017; Sadowski et al. 2017; Chael et al. 2018; Dexter et al. 2020a; Jiang et al. 2023) and M87 (Ryan et al. 2018; Mizuno et al. 2021; Dihingia et al. 2023).
@@
In this paper, we revisit the two-temperature GRMHD MAD simulations. The mentioned previous studies used several numerical codes assuming various torus sizes, magnetic field topologies, adiabatic indices, and grid resolutions to investigate the problem. It was not clear to us that the radiative output of these models is converged for MAD models with setups comparable to those in the EHT simulation library (Event Horizon Telescope Collaboration et al., 2022). The second motivation for the revision of two-temperature simulations is the fact that MAD models with $R~{}(\beta)$ prescription typically have too-variable total intensity compared to the observations of Sgr A* (Event Horizon Telescope Collaboration et al., 2022; Wielgus et al., 2022a). The origin of the discrepancy between the models and the observations is unknown. One possibility is that the $R~{}(\beta)$ model, which was initially developed for SANE simulations, is not suitable for MADs, and a more sophisticated electron model should be adopted. One could also ask how physically motivated the $R~{}(\beta)$ model is for electron temperatures when considering MADs. The latter is relevant for all sub-Eddington-accreting black hole systems. Finally, the measurement of black hole spin in EHT and other sources is typically accretion-model-dependent. A robust understanding of the dissipation processes in accreting plasma near the event horizon is critical for reliable spin estimates.
@@
In this work, we investigate how the numerical model parameters impact the radiative (total intensity and polarimetric) characteristics of the two-temperature MAD simulations. We first demonstrate that they are independent of the grid resolution used in GRMHD simulations. Next, we evolve models for longer times and compare the two-temperature models with $R~{}(\beta)$ models. Finally, having checked that the results are weakly dependent on the exact shape of the electron distribution function and that they remain unchanged even when radiative cooling is introduced (scaling models to the Sgr A* system), we compare the prograde and retrograde two-temperature MAD models to the newest multifrequency polarimetric millimeter observational data of Sgr A*. In contrast to most of the previous two-temperature simulations, in our comparisons, we focus on variability and polarization, which together are significantly more informative than the time-averaged or total intensity comparisons. We discuss the future prospects for investigating the thermodynamics of the GRMHD simulations.
@@
The paper is structured as follows. In Section 2 we describe our 3D two-temperature GRMHD simulations setups and outline the details of radiative transfer modeling. In Section 3, we report the results and compare the models with selected observations of Sgr A*. We discuss the results and conclude in Section 4.
@@ 2.1 GRMHD simulations
MAD simulations are carried by means of ebhlight, a radiative GRMHD code developed and made public by Ryan et al. (2015). ebhlight is a relativistic, second-order, conservative, constrained-transport code for stationary spacetimes. While the fluid part of the code is based on the harm routines (Gammie et al., 2003), the radiation part is based on grmonty scheme (Dolence et al., 2009).
All simulations are set in geometrized units where the length scale and time scale are set by the black hole mass: ${\mathcal{L}}\equiv GM/c^{2}$ and ${\mathcal{T}}\equiv GM/c^{3}$ (assuming $G=c=1$ both time and distance are measured in units of mass, M). The simulations start from the Fishbone-Moncrief torus (Fishbone & Moncrief, 1976) at the equatorial plane of a Schwarzschild or Kerr black hole. The torus is described by two parameters: inner radius $r_{\rm in}=20$M and radius of the pressure maximum $r_{\rm max}=41$M. The torus is seeded with weak poloidal magnetic fields described by vector potential,
@@
where $r,\theta$ are the radius and polar angle in Kerr-Schild coordinates, parameter $r_{0}=400$M, $\rho$ is the plasma density. The initial magnetic field is renormalized so that plasma $\beta_{\rm max}=100$.
The simulations use ideal equation of state with constant adiabatic index of $\gamma_{ad}=13/9$; i.e., we assume that plasma is pure hydrogen with nonrelativistic protons and relativistic electrons for which adiabatic indices are $\gamma_{\rm p}=5/3$ and $\gamma_{\rm e}=4/3$, respectively.
All simulations are carried out in mixed modified Kerr-Schield logarithmic coordinates, where the resolution is focused on the equatorial plane and toward the central region close to the black hole horizon. The grid stretches from within the black hole event horizon until $r_{\rm out}=1000$M.
The grid refinement convergence test runs are integrated for $10,000$M assuming four grid resolutions listed in Table 1. Convergence runs are performed for black hole spin $a_{*}=0$. Later in the paper, we consider both prograde and retrograde fiducial models (also listed in Table 1) with the following spin values: $a_{*}$=-0.9375, -0.5, 0, 0.5, 0.9375. The prograde and zero-spin models are evolved until $30,000$M, and the retrograde models are evolved for a shorter time, only until $14,000$M. In all our simulations, the accretion flows reach steady state within $r=20$M. Fiducial models have grid resolutions ($N_{r},N_{\theta},N_{\phi}$) = (240, 120, 128) (for $a_{*}$ = -0.5, 0, 0.5) and ($N_{r},N_{\theta},N_{\phi}$) = (266, 120, 128) (for $a_{*}$ = -0.9375, 0.9375). All models above are run with radiative modules of ebhlight turned off. We carry out one test run with radiative effects turned on. The parameters of the exploratory radiative GRMHD (GRRMHD) model are shown in Table 1.
@@ 2.2 Radiative transfer
We measure the convergence of the models by studying their radiative properties. To predict synchrotron emission from the simulations, we postprocessed the GRMHD snapshots using the ray-tracing relativistic polarized radiative transfer code ipole (Mościbrodzka & Gammie, 2018). The radiative transfer calculations are typically carried out starting at later times of simulations when the accretion flow is relaxed from the initial conditions (starting times, $t_{\rm s}$, and final times, $t_{\rm f}$, of all radiative transfer postprocessing simulations are listed in Table 1). We scale all GRMHD simulations using the mass and distance of the Sgr A* black hole. The models density scale is set by unit ${\mathcal{M}}$, a standard mass or accretion rate $\dot{M}$ scaling factor (e.g., $\rho_{Sgr~{}A*}={\mathcal{M}}/{\mathcal{L}}^{3}\rho_{code}$). In all models $\mathcal{M}$ is set to produce an average Sgr A* total flux of 2-3 Jy observed at a frequency of 229 GHz (Wielgus et al., 2022a). Table 1 lists each model ${\mathcal{M}}$ and corresponding $\dot{M}$. Given the electron temperature model (see next subsection), the imaging code produces synchrotron emission maps at a desired frequency $\nu$ (here 86–229 GHz) and a viewing angle $i$. Initially, all maps are computed for a default viewing angle $i=160^{\circ}$ selected based on the recent rediscovery of a transient hot spot orbiting around Sgr A* (Wielgus et al. 2022b, see also viewing angle estimates by Yfantis et al. 2024) and then later for additional viewing angles $i=150^{\circ},130^{\circ},110^{\circ}$. Note that for the modeled source, lower viewing angles are favored by EHT observations that revealed a symmetric ring (Event Horizon Telescope Collaboration et al., 2022); hence we do not model emission for high viewing angles, $i\sim 90^{\circ}$. Assuming $i>90^{\circ}$ also guarantees the correct sign of the circular polarization for the assumed polarity of magnetic fields in the GRMHD models. Our model map resolution is $256\times 256$ pixels and the field of view = $120$M $\approx 600$ $\mu as$. A time series of maps/images with a cadence of $\Delta t=10$M is synthesized into light curves in Stokes ${\mathcal{I}},{\mathcal{Q}},{\mathcal{U}},{\mathcal{V}}$. Following Wielgus et al. 2022a the modulation index $M_{3}{}\equiv\sigma_{\Delta T}/\mu_{\Delta T}$ (ratio of standard deviation to mean flux calculated on time intervals $\Delta T$ = 3 hr) is used to characterize light-curve variability in Stokes ${\mathcal{I}}$ at two frequencies, 86 and 229 GHz. We also calculate the spectral index in total intensity ($\alpha_{I}$ between 213 and 229 GHz), (Faraday) rotation measure (RM; between 213 and 229 GHz), and linear (LP) and circular (CP) fractional polarizations at 86 and 229 GHz. These particular quantities are studied because they are later compared directly to observations. In our models Faraday rotation is caused by relativistic and subrelativistic electrons within 100M from the event horizon; these electrons constitute the so-called internal Faraday screen.
@@ 2.3 Electron distribution functions
Following the scheme of Ressler et al. (2015), the ebhlight code tracks total and electron entropies that are used to calculate the nonadiabatic (viscous) gas heating rates and evolve electron temperatures. The scheme requires the provision of a subgrid model for the proton-to-electron heating ratio $Q^{+}_{\rm p}/Q^{+}_{\rm e}$. In this work, we adopt the prescription for $Q^{+}_{\rm p}/Q^{+}_{\rm e}$ developed by Kawazura et al. (2019, henceforth model K), which approximates dissipation in a turbulent cascade with a functional form
@@
In this model, most of the dissipation goes to protons in high $\beta$ plasma regions, while electrons receive most of the heating in low $\beta$ regions. Qualitatively and quantitatively, the K model is very similar to the turbulent heating model of Howes (2010). Studying significantly different options for $Q^{+}_{\rm p}/Q^{+}_{\rm e}$, such as, for example, electron heating by reconnection (Rowan et al., 2017), is beyond the scope of the current paper but is discussed in Section 4.
In ideal GRMHD models studied here using a conservative code, the total viscous heating is produced by truncation errors at the numerical grid level. One may conclude that no matter what fraction of this heating goes to the electrons, this heating will be completely artificial. However, for the turbulent torus problem, the grid-scale dissipation is set by the large-scale turbulence in the problem (see Section 3.1 in Ressler et al. 2015 for a more detailed discussion). The numerical scheme for electron heating implemented in ebhlight has been carefully tested against some analytic problems by Ressler et al. (2015) and Ryan et al. (2017) as well as Sadowski et al. (2017), who, except for the mentioned modifications, follow the same idea when evolving electron temperatures alone. We are not going to repeat these tests here. In ebhlight the electron evolution also does not include Coulomb couplings, unless radiative transfer is activated (see Section 3.2.1). However, those are not important in the low-density gas considered here. In ebhlight the electron temperature tracking is passive; i.e. the electron pressure is not accounted for in the GRMHD equations. More importantly, the heating of the proton (nonadiabatic/viscous) is not taken into account in the GRMHD equations as well; instead, the internal energy of the gas is evaluated from the total energy using an inversion scheme implemented in the code assuming $\gamma_{ad}$. The proton temperatures are computed using the total internal energy $u$, the plasma density $\rho$, and $\gamma_{\rm p}=5/3$ ($\Theta_{\rm p}\equiv(\gamma_{\rm p}-1)u/\rho$).
The simulations are two-temperature, but to make a connection with the previous studies, we use the same GRMHD runs and compute images/light curves using a parametric electron temperature model $R~{}(\beta)$. In this model, proton temperatures are calculated from GRMHD quantities, and the electron temperature is found using the formula below:
@@
where parameters $R_{\rm low}$ and $R_{\rm high}$ are temperature ratios that describe the proton-to-electron temperature ratio in strongly (low plasma $\beta$) and weakly (high plasma $\beta$) magnetized regions, respectively. To be consistent with Event Horizon Telescope Collaboration et al. (2022, 2024), we calculate the light curves for $R_{\rm high}=1,10,40,160$ and $R_{\rm low}=1$. The dimensionless electron temperature that is passed to the synchrotron transfer coefficients in ipole is calculated from
@@
with internal energy $u$ and rest-mass density $\rho$ provided by GRMHD model. The electron model assumes $\gamma_{\rm e}=4/3$ and $\gamma_{\rm p}=5/3$, self-consistent with GRMHD simulations in which electrons are relativistic and protons are nonrelativistic.
Models with the Kawazura et al. (2019) prescription for $Q^{+}_{\rm p}/Q^{+}_{\rm e}$ are tagged K, and models with parametric electron temperatures are labeled R1-R160 (e.g., model with $R_{\rm low}=1$ and $R_{\rm high}=160$ is denoted R160). The latter models are referred to as $R~{}(\beta)$ models.
It is reasonable to assume that the energy from the turbulent heating is used to form a nonthermal rather than purely thermal electron distribution function. We consider an exploratory nonthermal model using the $f_{\kappa}(\gamma,\kappa,w)$ distribution function (Xiao, 2006):
$f_{\kappa}(\gamma,\kappa,w)=N\gamma\sqrt{\gamma^{2}-1}\left(1+\frac{\gamma-1}{\kappa w}\right)^{-(\kappa+1)}$ (5)
where $\kappa$ and $w$ are parameters and $N$ is the normalization constant. The distribution parameter
$w=\frac{(\kappa-3)}{\kappa}\Theta_{e},$ (6)
where $\Theta_{e}$ is calculated form the electron entropy in exactly the same manner as in purely thermal K models (see Ressler et al. 2015). Equation 6 states that the entire turbulent energy is distributed into the $\kappa$ function. The parameter $\kappa$ is globally constant. Inspired by the solar wind studies, we assume $\kappa=4.25$ (e.g., Livadiotis et al., 2018). In ray-tracing, the radiative transfer coefficients for the $\kappa$ distribution function are adopted from Mościbrodzka & Gammie (2024). No cutoff is applied to the $\kappa$ distribution function at high energies because it makes a negligible difference when modeling (sub)millimeter emission.
All radiative transfer simulations are carried out only in regions where $\sigma<\sigma_{\rm cut}=1$ (where magnetization parameter $\sigma\equiv 2P_{\rm mag}/\rho c^{2}$). This is a standard procedure to avoid modeling emission from near-vacuum jet regions in the GRMHD model where the proton and electron temperatures are inaccurate (e.g., Event Horizon Telescope Collaboration et al., 2022). However, the millimeter emission from regions with $\sigma>1$ is small, as already shown by Dexter et al. (2020a, see their Appendix D).